OpenStudy (anonymous):

Use the properties of logarithms to evaluate log7343.

OpenStudy (loser66):

log base 10? I don't know how to do since 7343 = 7*1049 and both them are primes. I cannot get the answer without calculator.

OpenStudy (anonymous):

okay? thanks i guess :)

OpenStudy (anonymous):

a= 3 b= 2/7 c= 4 d= 1/7

OpenStudy (mathstudent55):

7^3 = 343 Is this problem log(base 7) of 343? If so, log(base7) 7^3 = 3*log(base 7) 7 = 3 * 1 = 3

OpenStudy (mathstudent55):

If the problem really is log(base 7) of 343, then you could write it as log7(343). When you write just log7343, it looks like you want log of 7343, which, without specifying a base, can be thought of log base 10 of 7343.